BUDDHIST PHENOMENOLOGY

Henk Barendregt

University of Nijmegen Toernooiveld 1 6525 ED Nijmegen The Netherlands

Proceedings of the Conference on Topics and Perspectives of Contemporary Logic and Philosophy of Science, Cesena, Italy, January 7-10, 1987, (Ed. M. dalla Chiara), Clueb, Bologna, 1988, 37-55.

Contents

1 The aim 3 1.1 Overcoming suffering . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 The small and large vehicle . . . . . . . . . . . . . . . . . . . . . 4 1.3 Detachment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.4 Insight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.5 The fruits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 Methodology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.7 Explaining apparent contradictions . . . . . . . . . . . . . . . . . 7

2 The path 8 2.1 The laboratory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2 The basic exercises . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3 The hindrances . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.4 Using the hindrances . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.5 Confronting suffering . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.6 Super consciousness . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.7 Disgust . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.8 Emergency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.9 Back in the world . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3 Again 15 3.1 Repeating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.2 The five groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.3 Sublime consciousness . . . . . . . . . . . . . . . . . . . . . . . . 18 3.4 Broken consciousness . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.5 Gluing consciousness . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.6 Types of consciousness . . . . . . . . . . . . . . . . . . . . . . . . 20

Abstract

Analytical philosophers have criticized some phenomenological texts from buddhism, existentialism and mysticism, because of the presence of logical contradictions. Being interested in those phenomenological texts, people with a different philosophical inclination sometimes make the following claims. “There are two ways of viewing the world: the rational and the irrational. The first view gives rise to science (with all of its drawbacks), the second one to the ‘higher’ truths of mysticism.” In his book Exploring mysticism (Penguin, 1975), F. Staal disagrees with such claims. He put forward the following views. 1. Mysticism consists of experience and is as such neither rational nor irrational. 2. As phenomenon mysticism can be studied in a rational way. 3. In order to do this, it is advisable to practice meditation in order to have first hand experience.

Following Staal’s program, this paper will first explain briefly how contradictions occurring in phenomenological texts can be understood. Then the main part of the paper will be devoted to a description of phenomenological data collected during the practice of buddhist meditation.

Acknowledgments

The information contained in this paper is obtained by following several practice periods of intensive meditation under the skilled guidance of The Most Venerable M.T. Mettaviharee, teacher of buddhist meditation in Amsterdam. I wish to thank Dr. Dick de Jongh of the University of Amsterdam, for improving the English text.

1. The aim

Buddhism has been described in many ways. It has been called a religion, a philosophy, a way of life and a psychology. Each of these descriptions is correct and emphasizes different aspects. Buddhism is a religion as it deals with questions of life and death and the meaning of life. It is a philosophy as it does not need the hypothesis of a god or of life after death. Buddhism is called a way of life as it teaches the way of non-violence and compassion. Finally, it is called a psychology as it investigates the working of our mind and distinguishes different types of consciousness.

There is yet another way to describe buddhism, namely as a course. The aim of this course is first the lessening and in the end the elimination of human suffering. This goal is to be reached by using experimental phenomenology: investigation of our consciousness as it is presented to us by means of self-observation. The method, both in theory and in practice, and some methodological considerations will be treated in this paper.

1.1. Overcoming suffering

If we find ourselves in a situation that is not agreeable to us, there are essentially two different ways to attempt to overcome the resulting suffering. On the one hand we can try to change the situation itself. On the other hand we can try to change our consciousness of the situation in such a way that it is no longer experienced as unpleasant. Depending on the circumstances and our possibilities, these attempts will be more or less successful.

It is clear that on the one hand science, technology and democracy have contributed considerably to the way our world can be controlled. On the other hand, we have buddhism, in particular the ‘theravada’ school (literary the ‘teachings of the elder’) that is focussed on the second possibility: eliminating suffering by changing our consciousness. Having completed successfully the course described above implies that one is accomplished in being able to change freely one’s consciousness. It may take many years, but the aim is worthwhile.

1.2. The small and large vehicle

At one time in history the theravada buddhists found themselves in a world having poor social-economic conditions. They were then criticized as “thinking only about their own enlightenment, while the world is burning”. This criticism was taken seriously. As a result so called ‘Mahayana’ (‘large vehicle’) buddhism developed. The followers of this branch of buddhism worked on improving both their consciousness and the world. To them it did not matter that the goal of overcoming suffering was going to be reached more slowly in this way: “we should reach nirvana as the last ones” became the bodhisattva vow of Mahayana buddhism.

The Mahayana buddhists used the pejorative term ‘Hinayana’ (‘small vehicle’) for the traditional theravada buddhism. The difference is, however, a matter of emphasis. A follower of theravada, too, will find it important to care for our environment. But during the theravada buddhist course the attention is focussed on changing consciousness. The sooner one has reached the goal the better, because with an improved state of mind, it is so much easier to create order around ourselves.

Mahayana buddhism became successful in China, Tibet, Korea and Japan, whereas the theravada school throve in Thailand, Birma and Sri Lanka. It is tempting to explain this phenomenon by the difference in climate in the countries involved. In the north living conditions are more difficult, therefore controlling the environment is more important and Mahayana developed.

As modern society is quite proficient in manipulating the environment, whereas the method of dealing with consciousness is little known, I expect in the future more influence of theravada buddhism. This paper is about theory and practice as taught in this lineage.

1.3. Detachment

One may wonder whether it is possible at all to overcome suffering by changing consciousness. And if it is possible, whether it is desirable to do so. We shall only discuss the first question. The second one is left to the reader.

Suppose we are in a restaurant and the waiter does not come. Even if we have plenty of time, we usually get somewhat irritated, often without noticing it. It may happen that we suddenly realize: “O, there is irritation in me”, and at once that irritation becomes much weaker or even disappears completely. This phenomenon is called ‘néantiser’ by Sartre. It feels like ‘melting away’ the emotion. It should not be confused with suppressing the emotion.

This is the key to the end of suffering: we do not view ourselves as somebody that is irritated, but as somebody that sees irritation. We do no longer identify with the emotion.

For a relatively light emotion as described, it is reasonably easy to apply the required detachment. This kind of phenomenon, of melting a light emotion, is relatively well-known. For stronger forms of irritation or of suffering in general, this method of objectification is more difficult to use. A strong form of concentration is needed in order to do this. Otherwise the irritation, for example, may immediately return after we have objectified it and consequently it keeps control of our consciousness. However, concentration, albeit very essential, will not be the most important tool in reaching complete detachment.

1.4. Insight

In order to follow the buddhist course one has to develop insight in the way our mind works. One of the fundamental statements of buddhism is that our consciousness is selfless. Our feeling of ‘self’ is seen to be a form of attachment that has to be overcome in order to eliminate suffering. It is remarkable that in order to obtain this insight one has to overcome a strong emotional resistance, the so called attachment to self. Easier to understand forms of insight will be described in sections 2 and 3. For example that we are manipulated by our emotions because these consist of several components that diligently reinforce each other.

1.5. The fruits

Having completed the buddhist path, or part of it, provides us with several sorts of profit. We have already mentioned insight in the functioning of our mind. Another effect is a considerable sharpening of our senses (something that can be obtained also through other kinds of training). Also we acquire the capacity to concentrate quickly.

However, the most important result, apart from the elimination of suffering, is freedom. Buddha said As the sea is pervaded by one taste, namely that of salt, in the same way my teaching is pervaded by one idea, that of freedom.

Some explanation is due here. Both Spinoza and Freud have remarked that our behavior is determined only partially by our conscious will but much more by something else. For Spinoza this something is our feeling, for Freud it is our unconscious. Often these distinct forces (conscious will versus feeling/unconscious) cause conflicts. Buddhism holds that if we are no longer attached to our feeling, then we are free.

1.6. Methodology

Phenomenology investigates consciousness as it is presented to us, we may say internally. As was mentioned above, progress in the buddhist course is correlated to gaining insight in the functioning of our mind. In fact it is necessary to have this phenomenological information in order to proceed on ‘the path’.

One may question the value of information obtained by self observation: “Is it possible to be objective and precise?” The answer is affirmative. The reason is that buddhist phenomenology derives its force from the systematic development of an essential discipline, namely meditation. During meditation mental tools are being built that are used for experiments. Experimental phenomenology is the principal method of meditation.

Nevertheless everything is internal, personal: with one’s own means one studies the working of one’s own consciousness. Therefore the objectivity of buddhist phenomenology may be questioned. The only thing that really convinces is taking part in the meditation experience. But perhaps the following comparison is reassuring. Mathematics is commonly recognized as a discipline of great precision and objectivity. Yet that subject depends essentially on a notion of truth that is internal (‘intuition’). Although in mathematics one uses calculations and formal proofs, there is always a need to have an intuitive judgment that verifies the correctness of the computation or the proof. Thus Husserl and Gödel rightly see mathematics as a part of phenomenology. Even if mathematics is about structures that are beyond us, for example about the set of natural numbers 0, 1, 2, ... , it is nevertheless the case that our contact with these structures runs via internal processes. Because of this phenomenological contact we are able to develop an objective mathematics concerning these natural numbers.

It is useful to compare buddhist phenomenology not only with mathematics, but also with a science, say physics. In physics the theory is built on experimental data. It is not true, however, that in order to study physics, one first learns all the possible experiments and then the development of the theory. On the contrary, experiments and theory are essentially intertwined. At first some simple theory is based on some empirical data. Using this theory new questions may be asked and new and more refined or totally different experiments may be performed. Then the theory is extended. Etcetera.

The situation is analogous with buddhist phenomenology. To the beginning student it is advised to have some confidence in the theory, but not to have blind faith. When experience is obtained, more and more of the theory can be verified. In this respect the theory is often compared to a map of an unknown country. When during a trip in that country the data on the map turn out to be true, only then we start to have real confidence in it. In chapter 2 and 3 the theory and practice of buddhist phenomenology will be described simultaneously.

1.7. Explaining apparent contradictions

Now we will explain how contradictions, which happen to occur in some buddhist texts, are possible. Suppose some part of reality U is described using some language L. Some of the regularities observed in L are in fact physical laws, but may be confused with logical laws. If we extend the reality U to U+, but keep as the describing language L, then statements may result that contradict statements made about U. Although the contradictions are only apparent, because the statements are about different ‘worlds’, it may seem that logical laws are violated.

An example will be helpful. Consider a tribe living on an isolated island. Vision of the tribesman is such that they can only see the colors black and white. In their description of the world they say: “Something is either black or white.” Although we know that this is for them in fact an empirical law, the people of the island are tempted to consider this as a logical law. Sometimes they use the words ‘white’ and ‘non-black’ interchangeably. On some day someone has a mystical experience. In our language we can say that that person has seen the color green. In the language of the tribe she says: “I have seen something very impressive. It was neither black nor white.” For most of the people of her tribe she was saying: “It was neither black, nor non-black.” Therefore on the island one may think she is speaking nonsense. However, we know that she is not.

There are, however, stronger contradictions. In his book Exploring Mysticism already mentioned, F. Staal discusses the following so called ‘tetra lemma’ occurring in buddhist texts.

It is not A;

it is not non-A;

it is both A and non-A;

and it is neither A nor non-A.

Even this contradiction may be explained. Simply consider again the tribe seeing only black and white. But now our mystic sees the color gray. Indeed gray is not white, not black. And it can be said that gray is both white and black. But also that it is neither white nor black.

I hope that the examples show that contradictions occurring in texts of mystics are not a sign that something essential is wrong. Nevertheless it is preferable that descriptions of altered states of consciousness are free from contradictions in the sense of logic. I will try to fulfill this requirement in sections 2 and 3.

2. The path

In section 1 we have already indicated briefly what is the key to the elimination of suffering. For example, by seeing in the right way that we are irritated, this irritation diminishes. This mental act of seeing what happens to us, is called mindfulness. The path, as the buddhist course is often called, consists of learning to become proficient in applying mindfulness. Once one has successfully completed the course, mindfulness is total and without effort. Then the mental formation of suffering can be fully seen and one is not touched by it anymore. This does not imply that one becomes cold or indifferent.

In order to get a better understanding on the nature of the path, it should be compared to a course to learn to swim. One can hear, read and speak about swimming, but progress can be made only by going into the water and doing one’s exercises. Something similar holds for the buddhist path.

2.1. The laboratory

The beginning student of the buddhist course is advised to practice meditation daily, for example for a period of one hour. However, essential progress is usually only made possible by so called intensive meditation. ‘Intensive’ means more than eight hours a day; it means engaging oneself every moment of the day with the act of mindfulness. By meditating intensively for a period of ten days, one may make definite progress.

It is clear that for complete devotion to mindfulness one has to live for some time under special conditions that are suitable for ‘the work’. Traditionally this happens in the monastery, the ‘laboratory’ for the mind. The word ‘laboratory’ may sound peculiar to some readers. Others may object to ‘monastery’. What is important are the following conditions.

• Right living. A private room in which one is not disturbed.

• Right temperature. Not too cold, not too hot.

• Right food. Not too much, not too little; sufficiently nutritious.

• Right friends. Persons that take care of preparing food for the students should respect the meditation work that is being done. Moreover an experienced teacher is essential.

During the intensive meditation period (retreat) one does not speak, except with the teacher in a daily interview in which he follows the progress of the student and gives hints.

The remainder of the paper is a description of the exercises and the resulting experience as obtained during some meditation retreats. Although these experiences are my personal ones, the general tendency can be expected by anyone who follows the method.

2.2. The basic exercises

There are three basic exercises: walking, sitting and prostrating (bowing) all with the aim of increasing mindfulness. These exercises have to be performed with the utmost care. To outsiders (and many students of buddhism as well) they will seem absurd.

Walking. One starts by standing with the feet next to each other. One says slowly in oneself: “standing, standing, standing”. Then very slowly the right foot is lifted and put forward one step. Simultaneously one thinks: “Right goes thus”, the thought exactly following the movement. After this step there is a little pause. Then one makes a step with the left foot and the thought follows: “left goes thus”. One continues like this for one hour. The next time one walks, the mindfulness may be increased by naming more carefully the movements of the feet. First “lifting, treading”; “lifting” exactly when the foot leaves the floor and “treading” exactly when the foot touches the floor again. Then one uses consecutively (each phrase for naming the movements is used for at least one hour): “lifting, going, treading”; “heel up, lifting, going, treading”; “heel up, lifting, going, dropping, treading”; and finally “heel up (extremely), lifting, going, dropping, treading, pressing front (foot), pressing back (foot)”.

Sitting. One sits on the floor on a cushion with one’s back straight and crosslegged. For example in full or half lotus position. The eyes are usually closed. One does not move, even if there is pain in the legs, one has an itch or something the like. One pays attention to one’s breathing by following the movements of the abdomen: “raising, falling” or “raising, rest, falling”. Again the naming follows the body. Sitting is done also for the period of an hour.

Prostrating. One starts by kneeling. One put one’s hands together with the fingers stretched upwardly. In this position the hands are put before the breast and one thinks ‘purity’. Then the hands, still in the same position are put before the forehead and one thinks ‘compassion’. Finally one bows deeply, the elbows touching the knees and the lower arms and hands on the floor while keeping one’s legs in a straight line. One puts the forehead in front of the hands on the floor, in this way raising one’s behind. One thinks ‘wisdom’. This prostration exercise is done each time for twenty minutes.

It is important to consider the basic exercises not as a walking, sitting or bowing exercise, but as training in mindfulness.

During the retreat only these three exercises are done, one after the other non-stop. The only interruptions consist of the meals, contact with the teacher, practical actions (like washing oneself and dressing) and sleeping. It is important to keep the meditation in the form of mindfulness also during the meals and the practical actions. One continuously pays attention to everything that is there, to the input of our senses.

2.3. The hindrances

At this moment things are becoming interesting. It turns out to be impossible to perform the basic exercises correctly. In ordinary life we are used to avoid discomfort as much as possible. If our position is uncomfortable, then we move. During meditation this is not done. Therefore all kinds of problems arise as disturbances. These prevent the student from paying attention to the meditation object (like our footsteps).

There are five groups of ‘hindrances’: lust, hatred, drowsiness, doubt and restlessness. Each of these will be explained.

Lust. By this we mean feelings (or thoughts) that we like and cause desire. For example, the student may be thinking about the coming meal instead of paying attention to his footsteps or breathing. Lust may arise in many different ways. A characteristic attribute is that one is manipulated by it and cannot escape its grasp.

Hatred. This refers to all unpleasant feelings that cause aversion. This emotion manipulates us too. During sitting the student may be in pain, often extremely so. Some part of the body may itch. Or the student may be irritated. Because it is a rule not to move, the student is forced to confront suffering.

Drowsiness. This is a well-known phenomenon. After a period of being bored, the student may rather fall asleep, than pay attention. Sometimes, during sitting, this actually happens. The attentive student will notice that he gets drowsy not because he is tired, but because he has to do something that he does not like and that causes mental and physical resistance.

Doubt. The continuous alternation of pain, boredom and desire for something else makes the student wonder what he is doing. He is doubting whether he should not have spent his ten days on a nice holiday. The reasons for coming to the monastery are questioned.

Restlessness. Also this word is self-explanatory. Although it is often a negative experience, it is not classified as hatred. The reason is that restlessness may also be accompanied by lust. As an example in daily life we know the awaiting a meeting with a new love. In meditation restlessness is considered as a hindrance, because it prevents mindfulness.

2.4. Using the hindrances

At first the student will try to fight the hindrances. Discipline and concentration are said to be important. But even when trying very hard, it turns out impossible to overcome the disturbances. Disappointed and perhaps desperate the student consults the teacher.

An important method to continue practice, in spite of the disturbances, consists of so called ‘switching’. This will turn the hindrances into something positive. The method is as follows. As soon as a hindrance arises in the consciousness of the student, he tries to push it away in a gentle but firm fashion. If this succeeds, he then returns immediately to the original object of meditation, like the breathing. In case it is not possible to avoid the influence of the disturbance, the student changes his object of meditation. He switches to the hindrance itself and names it as ‘irritation’, ‘lust’ or the name of one of the other hindrances. As soon as the hindrance stops being disturbing, the student switches back immediately to the original object of meditation.

In this way it is almost always possible to exercise in mindfulness. An exception is the hindrance of drowsiness. If this occurs, the student may start the walking or even better the prostration exercise. In this way he does not fall asleep and moreover, obtains new energy.

Using the method of switching the student will understand better what genuine mindfulness is. This is not always obvious. It may happen that the student thinks he is mindful, but in fact is not fully so: there is one thing in which the student is more involved, namely concentration. By switching diligently between concentration and the object of meditation, the student will notice the difference between mindfulness and concentration.

2.5. Confronting suffering

Even if the method of switching helps the student with his meditation, it will not always be pleasant. On the contrary, it is almost certain that after skillful switching the hindrances will grow stronger and start alternating rapidly. The resulting hardship will be extreme. Strong pain or anxiety; restlessness.

At this point the student starts complaining to the teacher. He wants to discontinue the retreat. For him meditation has become like dying. However, the teacher advises the student to continue. “In this life we often ‘die’. Each time when a new desire starts we are ‘born’. If the desire is strong, then we ‘burn’. Finally if it turns out to be impossible to obtain what we want, then we ‘die’. It is not that terrible. In our meditation we are working towards cutting through this circle of birth and death, this so called ‘samsara’. Finally we will not be born anymore and then we are free.”

The student who has managed to endure the retreat so far will usually have enough discipline to continue. He names his suffering: “This is suffering; this is suffering.” But as the feeling is very strong, it does not melt away. The student is hoping that this will happen soon.

2.6. Super consciousness

Exactly this hope, this desire causes the discomfort to be even more strong. The situation is constantly deteriorating. Finally the student gives up all hope for improvement. Even then there is extreme suffering.

At that point something unexpected happens. Suffering remains, but suddenly it is not bothersome anymore. The student clearly sees the following. “This is my body, this is my mind; both are burning, but it does not matter.” From that point on the method of mindfulness also works for strong forms of suffering. This probably has been made possible as the student has developed a strong form of concentration.

This all may sound unbelievable. Perhaps the following may relieve some doubts. Suppose one pronounces repeatedly a particular word, for example the word ‘rubber’, then it may seem that the word looses its meaning. Or rather the sound ‘rubber’ and the concept ‘rubber’ are being separated. Usually these two components are melded together in our consciousness. This phenomenon is well-known and is called a ‘jamais vu’ or a ‘semantic fatigue’. Something similar may also happen with our suffering. Pain is being decomposed into two components: ‘pain as pure feeling’ and ‘pain as something to run away from’. After this decomposition, using mindfulness, it is relatively easy to melt away the second part.

This method of splitting consciousness into various parts is characteristic for meditation as taught in theravada buddhism, the so called vipassana meditation. As a result one is able to understand better the functioning of consciousness. The word ‘vipassana’ means ‘insight’.

The consciousness of the student is now getting more and more ‘pure’. The hindrances are less and less frequent. And if they do happen to come, it is easy to use mindfulness in order to melt them away. The student is able to see the growing of every hindrance; and then he uses mindfulness “as a gun to shoot it away”. This is how it feels at least.

If the student is a ‘good boy’ and applies mindfulness also to feelings of lust, then a strong and stable consciousness will result. All phenomena are the same as before and all mental faculties are the same. But there are almost no disturbances anymore. Pain is strong, but far away.

The hindrance of doubt seems to be conquered completely. “It is true indeed that it is possible to improve one’s consciousness by means of meditation” the student is thinking. [As the student would find out later, a remarkable aspect of the new consciousness is that some of one’s ordinary skills are impaired. Writing is rather difficult, as if one is doing it with the wrong hand. Reading is extremely tiresome. Estimating distances does not work well either. Objects seem nearer than they really are. Sometimes one is not well-balanced while walking. After some trial and error one quickly is able to re- establish correct estimation of distances. The left-right aspects of the walking exercises play an important role in repairing the disturbances in one’s equilibrium.]

The altered state of consciousness satisfies descriptions occurring in mysticism. The student is taken by rapture and feels omnipotent. By melting away these kind of narcissistic feelings the consciousness of the student is getting to be even more stable. It satisfies the statement of the Hindi

Atman is Brahman

(the self is god). The student thinks he is enlightened. However, as will turn out soon, there is still a long way to go.

2.7. Disgust

The student is now continuing to meditate having a consciousness that is neither pleasant nor unpleasant, but is nevertheless very exalted. There is almost nothing to do. The hindrances are gone and being mindful on for example breathing is effortless.

At one moment the student notices that it becomes difficult to keep his super consciousness. More and more pressure is growing on his mind (or is it the head?). Even if it does not bother him, it is a threatening omen.

At any price the student wants to keep his exalted consciousness. But most unfortunately he cannot keep his strong concentration any longer and he falls back on what used to be his old consciousness. It turns out that this consciousness is totally neglected. It looks like a giant abyss. Desperately the student tries to regain his exalted consciousness. It does not work. Then suddenly the student looses all grip and falls in the abyss. The experience is indescribable. The following symptoms are some indications of what happens.

• Chaos: perception is completely confused; everything is turning.

• Anxiety: there is an ‘infinitely’ strong anxiety; all other anxieties seem to be derived from this basic one.

• Disgust: one is extremely sick; the stomach turns and one has to vomit.

• Craziness: it feels as if one is crazy; the mind and the body are present, but there is no ego anymore that controls everything.

The experience is more terrible than death, at least that is how it feels like. Nevertheless thinking operates as usual.

Panic stricken the student goes to the teacher, who is available 24 hours a day during the intensive meditation retreat, precisely for these cases of depersonalization. The student hopes that the teacher can perform some miracle. But that is not what happens.

The teacher reacts in an unexpected way. He says the following. “Remember the theory. Our existence has three characteristics: changeability, suffering and selflessness. Changeability means that no phenomenon is permanent. Even the visual image that we have of a solid object is not constant. If we look well, we will see many small fluctuations in the image as perceived. In this changing world we look for some hold, something that is constant. In our ignorance we consider our ego, our self, as a fixed remaining entity. But that is an illusion. Why would our ego feel the need to be consolidated, if it were really permanent? Well, the changeability causes the circumstances in which we live to be often in conflict with the ideas of the ego and friction results. That is suffering. Finally, selflessness means that the phenomena out of which our consciousness is built up, are essentially beyond our control. These three attributes of our existence, changeability, suffering and selflessness are in fact three aspects of one single truth. Now you have experienced yourself these three characteristics. This is important. Just continue your practice.”

But the student is not at all interested in meditation anymore. He just wants to stop his depersonalization. The teacher notices this and gives specific instructions. “Instead of prostrating for twenty minutes, you should do it from now on each time twenty minutes longer. That is, first twenty minutes, then forty, then sixty and so on. Work well on your mindfulness. Then the nausea will disappear by itself. However, if it comes back, do not name it as ‘nausea’ but as ‘knowing’. Do not feed it it.”

2.8. Emergency

Back in his room the student has lost all motivation and does not want to continue the practice. However, stopping for a rest deteriorates the situation: the abyss becomes more threatening and nauseating. In utter despair he follows the instructions of the teacher. After a couple of hours it turns out that the exercises have a calming effect. Especially the prostrations function well against the sickness. Then the student wants to relax. But each time he slackens his mindfulness, he is pushed back into the abyss. It is like having a knife on one’s throat forcing one to meditate fully. It is a matter of life or death. Being tired, the prostration exercises give new energy.

After having worked for many hours like this, being exhausted, the student is not sure whether he can go to bed. It may induce the nausea again. The student goes to the teacher and asks whether it is possible to go to sleep. “All right, sleep with mindfulness”, is the answer.

The student sleeps remarkably well. However, the next morning the emergency is still there. Only after some days of extremely hard work, the student re-establishes a stable consciousness without nausea. His ego feeling returns, albeit in another more light form. The student keeps some suspicion towards the method.

2.9. Back in the world

Returning from the laboratory the student enters ‘the world’. It turns out that he can function normally. Eventually this brings back some confidence to the student. However, sometimes there is nostalgia for the feeling of ego that he has left behind. But the new form of ego also has its advantages. An irritating situation can be relativized more easily. And indeed it turns out that all anxieties are corollaries to the fundamental one. By knowing this one, all other anxieties have become less serious.

3. Again

It may take a while before the student is motivated again to join an intensive meditation retreat. One of the possible motives is the memory of the super consciousness described in 2.6. Another reason may be, that on the one hand the student has gained confidence in the theory as parts of it turned out to be true, and on the other hand the theory promises much more. A description will be given now of the work done after one has chosen to continue the practice.

3.1. Repeating

The student notices that he has to start from scratch. The technique learned at the previous retreat has been lost. Fortunately progress is much faster now.

Again there are hindrances. Again they can be used to increase mindfulness. Again there is strong pain or strong suffering. Again there arises strong concentration, making the extreme pain easy to take. Again a stable consciousness results.

However, the teacher is not satisfied. “If there is strong pain, but you do not mind, then you first form the pain, while you throw it out later. That is a waste of the energy that it all takes. You should prevent the formation of the pain and in general of your feelings. Then you do not need to throw away anything. Cut through the addiction to your feeling. Stay with the input, with the contact of your senses.”

The student is not able to do that. He again develops a super consciousness. Again he notices that he cannot keep it up. When he falls this time, the nausea is less strong, as it is a known phenomenon by now. After only an hour the consciousness of the student is stable again.

3.2. The five groups

In order to proceed from here, the following facts from the theory should be known.

Our phenomenological awareness (as always both physical and mental) can be divided into five groups (‘skandha’ in Pali; usually translated with ‘aggregate’):

1. input (of our senses);

2. feeling;

3. perception;

4. output (conditioning, will);

5. consciousness (state of consciousness).

A more elaborate description will be given now.

1. Input. Everything that we are aware of enters through one of the six senses. These are characterized by the following verbs: to see, to hear, to taste, to smell, to feel (to perceive by touch) and to think. The first four senses, namely the eyes, the ear, the tongue and the nose do not need further description. The fifth sense, bodily feeling, has many manifestations. For example itch, pain and cold. There is no special location for this sense. The sixth sense is the mind. This is usually not considered as a separate sense, but rather as the place where all senses come together and are being processed. There are, however, some good arguments for classifying the mind as a sense. If we look up something in our memory, we obtain information that was not there before. If we have some thoughts after seeing some object, then, although that thinking is a consequence of the seeing, it is a separate input for our awareness. In the buddhist theory there is also processing of the input, but this is placed in the groups (2) - (5).

2. Perception. This group will be discussed first, because group (2) is somewhat more complicated. After the senses have delivered the input to us, perception makes recognizable units from the data. It makes a classification; as they say “it distinguishes this from that”.

3. Feeling. This group should not be confused with sense number five, the feeling of touch, itch, cold, etcetera. In group two we give a value judgment to our sense input. We say “This is pleasant, this is unpleasant and this is neutral.” Our feeling indicates whether we like to have a phenomenon in our consciousness or whether we don’t; or that we are indifferent to it. Indeed, according to the theory feeling comes in three forms: positive, negative and neutral. Because buddhism sees feeling as the direct cause of our behavior, group (2) is considered as very important. Positive feeling is equated with lust; negative feeling with hatred; and finally neutral feeling with ignorance. Lust, hatred and ignorance are seen as the three fundamental sins. Of these three ignorance is the most persistent one, since it is hidden and can change suddenly into lust or hatred. The reader may wonder why positive and negative feeling (‘desire and sadness’) are repudiated. One may reason as follows: “All right, one is better off not having negative feelings, but don’t the positive ones form the meaning of life?” To this one may reply that it is certainly the case that for example sex and good food are pleasant. But the desire for it has a hold on us, it manipulates us. The goal of buddhism is to be free from this desire. The mentioned delights will remain pleasant, even after one is enlightened. But then they are pleasant on the level of group (1), the sensory data. Although the following simile is not perfect, one may compare the enlightened enjoying- without-desire to listening to beautiful music. We enjoy the music of Mozart, but we do not cry when the symphony is over. Having ‘positive feeling’, in the technical sense of buddhism, means desire that is constantly directed to the object of desire, even if it is absent. A comparison with Calvinism is appropriate here. In this persuasion just as in buddhism lust is considered a sin. However, in Calvinism it is something that is given a negative value: it is something that is not allowed and one should be ashamed of it. In buddhism lust is considered as something that is present in us. It is considered a sin because it is the cause of suffering and bondage. But that is all.

4. Output (conditioning, volition). According to buddhism our behavior is conditioned. This applies to everything we do, both with our body and with our mind. Cybernetically this conditioning can be viewed as the output of our body-mind complex: our nerves are firing to control our movements and our thoughts. Phenomenologically conditioning is correlated to volition: we usually have the impression that we want to do whatever we are doing. However, this volition can be seen as a side effect. Therefore it is adequate to describe group (4) simply as output.

5. (State of) consciousness. In the first place consciousness is the integrating factor that unifies our input and its processing. So there is a difference between input, perception and consciousness. Input as such is still disconnected. (Sometimes when it is dark we may see that our visual input consists of small fluctuating parts.) Then, perception classifies the input, but only locally so. Finally consciousness gives a global image.

In the second place group (5) refers to our mood, to our mental state. It will turn out that the way our consciousness unifies our awareness is strongly dependent on our feeling.

There is a strong interaction between the five groups. Feeling results from input, but in a manner that depends on perception and consciousness. If someone says something rude to us, then the resulting feeling depends on who has spoken to us, that is, depends on perception. Moreover the result of the unfriendly words depends on our mood of the moment. That is, our state of consciousness (group (5)) also influences our feeling.

Conversely, feeling plays a central role in perception, conditioning and consciousness. In the next sections we will give evidence for this. Because of its central role, feeling is therefore placed as second group, directly after input.

The views of Spinoza and Freud concerning the bondage of our behavior are related to this theory. According to Spinoza our feeling determines our actions. According to Freud the determining factor is our unconscious. Buddhism reconciles these two views. Also according to this theory feeling determines our actions; however, large parts of that feeling are unknown. That is another reason why our ignorance is considered to be a sin: it determines our actions without our realizing it. And in addition we usually make rationalizations about our motives for those actions.

3.3. Sublime consciousness

The student analyses his experiences in the light of the theory of the five groups. Using strong concentration, he is able to endure strong pain, without being bothered by it. Probably that is because the connection between feeling and conditioning is interrupted. A drawback is, that it takes a considerable amount of energy. If one could disconnect the link between input and feeling, then perhaps suffering could be overcome much more easily. For this it is necessary to know exactly how feeling is formed.

Work is resumed. As usual the pain slowly develops during sitting. As soon as the pain becomes unbearable the student automatically switches to super consciousness: “Ouch, now I can take it.” But the teacher is dissatisfied: “You should not hide your pain, but stop forming it!”

The student is increasing his mindfulness. As soon as a little pain arises, he sees it. But he also notices, that directly after the pain is melted away using mindfulness, a whiff of pleasure arises because the pain is gone. Previously he had failed to notice this.

It now becomes an exercise in swiftness of the factor mindfulness. It runs as follows. Pain; melting; pleasure; melting; pain; melting; pleasure; melting; etcetera. At first the student is not well synchronized with this flux of phenomena. Then it goes better. Suddenly the student gains momentum. Part of consciousness is disappearing. Almost the student becomes nauseated again. But using mindfulness the formation of the sickness is prevented. And then, quite unexpectedly, he can do it: siting quietly for a long period without having pain. It simply is not being formed.

Now things are going quite well and effortless. Consciousness is becoming calm and stable. Because little effort is required much more calm as before. The consciousness described in 2.6 is now seen quite differently. As compared to the present consciousness that previous super consciousness is thick and treacly, due to the presence of neutral feeling. At the time it seemed to be the utmost experience.

The student enjoys his present sublime consciousness. It is stable. Having had previous experience with an altered state of consciousness the student wonders whether he can loose it again. At first sight there seems nothing to loose.

3.4. Broken consciousness

Suddenly something happens. There is new pain. The student is unable to prevent its formation. How come? The pain-pleasure method does not work. Moreover it is not clear whether or not the student is bothered by the pain. But it is quite strong.

The teacher is consulted. He says: “The pain (and more generally the feeling) is supported by your perception and consciousness.” The student does not understand this remark. The teacher continues: “While working with the pain, you may try to notice small fluxes in your consciousness.” The student asks: “Should I name this and make it melt away using mindfulness?” The teacher: “Yes, but only do this when you have confidence in your practice. You will interrupt the continuity of your consciousness, of your existence. Everything will start turning like being in a roller coaster. Then continue your practice, even if you feel sick. After a while-no one can say how long it will take-the turning will stop. Do not try to step out of the roller coaster, otherwise you will fall down.”

The student, who by now has got used to strange phenomena, is quite intrigued by these words. He continues the practice. Walking. Sitting. Pain is not being formed. After long sitting pain is being formed. At one moment it becomes almost impossible to melt away the pain. Then comes a strong outburst of pain. An even stronger blast of mindfulness is able to melt away the pain. By properly paying attention the student notices that his perception-right after the melting away of the pain-disappears locally for a short moment. The vision of an object laying in front of the student melts away for less than a second. It leaves a hole in his perception. The student tries again. Pain; melting; hole in perception. Pain, hole, pain, hole, ... ; faster and faster.

Then suddenly pain is not formed anymore. But consciousness collapses also. There is only awareness, with every now and then a puff of perception or consciousness. Intellect appears to be functioning normally. The holes in consciousness are turning around quickly. During the nausea described in 2.7 this must have happened also. But then the student did not see any details because of lack of mindfulness. Also perception of the body is disturbed. During walking the room seems to be a jolting carriage. The student does not become sick, because mindfulness watches his stomach well.

The experience is not pleasant, but not terrible either. However, when after two hours everything still keeps turning, the student starts to worry. Then he develops restlessness and wants to have a stable consciousness. There is nevertheless no nausea nor anxiety.

3.5. Gluing consciousness

The student decides to consult the teacher. His room is at a different wing of the monastery. Walking there, while developing strong mindfulness, the consciousness of the student suddenly is glued together. It just happened, without doing anything special. It is a new consciousness; yet everything in it is the same as before. The student changes his mind and starts walking back to his room. Then his mindfulness weakens and he falls back to his broken consciousness. This makes him decide to go to the teacher after all.

After the student has told what had happened, the teacher says: “Your practice is going well. Especially the way your consciousness has been glued together. Feeling will be substituted by mindfulness. In case your consciousness breaks again, name it as ‘knowing’. Do not identify with either the new consciousness, nor the broken old consciousness.”

After the return to his room, it takes a couple of hours for the student’s consciousness to stabilize. Then he starts analyzing what has happened.

The functioning of consciousness is apparently analogous to that of a television screen, consisting of rapidly fluctuating images. Normally we have the impression of continuity. Probably the reason is that feeling works as a kind of glue for these separate fluctuating images. By the meditation technique feeling is (temporarily) switched off. Therefore there is no more glue and consciousness will show itself in its broken form. The student does not understand well how everything works, but the phenomena are interesting.

3.6. Types of consciousness

Back to work. The exercises are going well. There are few disturbances and sublime consciousness results. Pain is not being formed. However, later pain arises again. When this pain is being melted away, consciousness starts breaking again. The student thinks “O, that is all right; I know how to depersonalize!” But he is wrong, it is going to be tough. The depersonalization is different this time and as this was unexpected, the student is getting restless again. Working hard makes consciousness stable. It becomes sublime again, even more so than before.

After continuing the work, the student falls again in a depersonalization. He becomes angry: “There is no end to this!” This emotion of angriness has an unexpected effect. The broken consciousness is glued together instantaneously. But how ugly it is. Everything exhales malignity: the walls are angry, each footstep is angry, everything is angry.

Work is continued. Consciousness breaks and the angriness in the walls suddenly disappears. The student tries whether consciousness can be glued together also with other feelings. First he tries lust. It works; when consciousness is glued together with lust everything becomes luscious. The student has no objection and keeps this type of consciousness. In the dining room he notices, however, that he is taking too much food. A habit he had overcome already at the beginning of the retreat. When the desert comes he promises to himself not to take too much. Nevertheless, he is taking more than he can eat comfortably. Apparently his hands took more than his mind wanted.

Back in his room the student decides that a consciousness glued together with lust is undesirable and he gives it up. During the resulting depersonalization the student happens to feel insecure. This time, by accident, the broken consciousness is glued together with anxiety. Everything exhales this emotion. The student is walking around in the surrealistic world of a phobic.

After having made a more quiet consciousness, the student goes to the teacher for his daily interview. The teacher is dissatisfied about these experiments. “You have to continue the basic exercises. Feeling has to be substituted by mindfulness. Do not glue with feeling, but with mindfulness. The difference is that you are dependent on your feeling. Mindfulness on the other hand sets you free. Be not attached to particular forms of consciousness. Let them come and go by themselves. The final goal is freedom. When you will have it, sublime consciousness may arise as often as you want. However, if you make some type of consciousness now, then you want to do that. As a result there is hidden feeling in it and you will not be free. This means that for example your sublime consciousness will depend on something. Therefore it will be impossible to keep it. When finally you will be free, then you do not depend on anything, you have become ‘unconditioned’.”

The period of intensive training is over, however. With good feeling the student returns home. So the work is not finished yet.

Postscript

The way to the elimination of suffering is the way to eliminate attachment to self. This means the elimination of our addiction to feeling. Personal experience of a small part of the way is described here. Striking occurrences were the following.

• Strong pain, without minding it;

• super concentration and super consciousness;

• depersonalization (very strong withdrawal symptoms);

• sublime consciousness;

• stabilization either by means of feeling or by means of mindfulness.

The possibility of using mindfulness to relativize problems is commonly known, albeit that usually no terminology for this act is given. The most striking aspect of the method of vipassana meditation is that this mindfulness, if combined with effort and concentration, can be used in a very powerful way.

The path is not always pleasant, but it is quite varied and remarkable.